# Mitigating distribution shift in machine learning-augmented hybrid simulation

### v1

@@ 2.3 Learning the unresolved model and distribution shift
We assume that the resolved part $\mathcal{L}$ is known, or we at least have access to a gray box that can perform its evaluation and compute its gradients. While the unresolved part is unknown, we can access data tuples $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$ which are obtained either by physical experiments or accurate but expensive numerical simulations. The goal is hence to construct an approximate mapping $(\mathbf{u}_{k},t_{k})\mapsto\mathbf{y}_{k}$ from this dataset. Viewing this as a supervised learning task, there are many existing methods based on empirical risk minimization [19, 20, 24]. One of the simplest method is based on minimizing the $L_{2}$ loss function. Given a parameterized model $\phi_{\theta}$, either a classical one with unknown parameters or a surrogate one such as a neural network, we learn from data the optimal value of the parameter $\theta$, i.e. $\widehat{\theta}=\arg\min_{\theta}\mathbb{E}_{(\mathbf{u},\mathbf{y})}\left\|\mathbf{y}-\phi_{\theta}(\mathbf{u},t)\right\|^{2}$. However, we emphasize that the data assumption in learning of dynamical processes differs significantly with the traditional i.i.d. assumption in statistical learning [20]. As the data is composed by several simulation trajectories, those belonging to the same trajectory will have large correlation. Therefore, it is reasonable to expect that simple empirical risk minimization (which many existing MLHS works employ [46, 42]) may not be the optimal method. We will demonstrate this by using the least squares estimator as a baseline comparison to our proposed approach Section 4.
@@
After determining the parameters of the unresolved model, one can perform hybrid simulation to obtain new trajectories, i.e.
$\widehat{\mathbf{u}}_{k+1}=\widehat{\mathbf{u}}_{k}+\Delta t\mathcal{L}(\widehat{\mathbf{u}}_{k},\widehat{\mathbf{y}}_{k},t_{k}),\quad\widehat{\mathbf{y}}_{k}=\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k},t_{k}).$ (8)
@@
Notice we add hat superscript to all the quantities related to this simulated dynamics to distinguish them from the ground truth $(\mathbf{u}_{k},\mathbf{y}_{k})$. We measure the performance of the simulator by the error along the whole trajectories, i.e.
$\min\sum_{k=1}^{n}\left\|\mathbf{u}_{k}-\widehat{\mathbf{u}}_{k}\right\|^{2},$ (9)
@@
where $\mathbf{u}_{k}$s are the true trajectories and $\widehat{\mathbf{u}}_{k}$s are the simulated one with the same initial condition.
@@
We begin by illustrating the issue of distribution shift by analyzing of the trajectory error. For simplicity of presentation, we assume that the model hypothesis space $\{\phi_{\theta}:\theta\in\Theta\}$ is bias-free, meaning that there exists $\theta^{*}\in\Theta$ such that $\phi=\phi_{\theta^{*}}$. In the general case where the hypothesis space is universal but not closed, this equality would be replaced by an approximate one, but the argument follows analogously. Comparing Eq. 1 and Eq. 8, we have
$\ \widehat{\mathbf{u}}_{k+1}-\mathbf{u}_{k+1}$ (10)
$=$ $\ \widehat{\mathbf{u}}_{k}-\mathbf{u}_{k}+\Delta t\left(\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\theta_{*}}(\mathbf{u}_{k}))\right)$
$+\Delta t\left(\mathcal{L}(\widehat{\mathbf{u}}_{k},\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))\right).$
@@
The error introduced by moving one step forward is thus decomposed into two parts: the error associated with the estimator $\widehat{\theta}:\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\theta_{*}}(\mathbf{u}_{k}))$, and the other error $\mathcal{L}(\widehat{\mathbf{u}}_{k},\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k}))-\mathcal{L}(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))$. The former resembles the error appearing in classical statistical inference problem under the i.i.d. setting since $\mathbf{u}_{k}$ follows an identical distribution of the training dataset. However, the latter is different in the sense that $\widehat{\mathbf{u}}_{k}$ and $\mathbf{u}_{k}$ may come from distributions of their respective driving dynamics, which are different since $\theta_{*}\neq\widehat{\theta}$. This may be due to noisy labels, few data points, or having an under-determined system. This difference between MLHS and statistical learning is also emphasized in [51]. This latter error is rather akin to the issue of stability in numerical solution for partial differential equations: although the error in each step is relatively small, it can accumulate exponentially as time step iterates, driving $\widehat{\mathbf{u}}_{k}$ to completely different regime where we have not observed in the data $\mathbf{u}_{k},k=1,\cdots,N$. As a result, the data-driven model $\widehat{\phi}$ can no longer be trusted to be accurate, since it is trained on data from a different distribution. This may then lead to a vicious cycle, where further deterioration of trajectory error occurs, leading to increased discrepancy between the distributions. Consequently, one can understand the distribution shift as a large discrepancy between the true distribution of $\mathbf{u}_{k}$ and simulated distribution $\widehat{\mathbf{u}}_{k}$. This problems occur commonly in MLHS. For example, [28, 29, 53, 39] all observe unphysical solutions when substituting Reynolds stresses models calculated from direct numerical simulation database into RANS. Ref. [53] attributes this to the ill-conditioning of the RANS problems and propose a method to treat Reynolds stresses implicitly, while [30] uses coupling modes perspective to numerically study the stability and convergence of RANS and [18] derives the propagation error of RANS simulation error mathematically. However, a general theoretical framework to understand and combat distribution shift in MLHS is lacking. The current paper aims to make progress in this direction.
@@
Let us illustrate the problem of distribution shift using numerical examples. We solve the 2D FitzHugh-Nagumo reaction-diffusion equation over a periodic domain $[0,6.4]\times[0,6.4]$, whose dynamics is given by
$\frac{\partial\mathbf{u}}{\partial t}$ $=D\Delta\mathbf{u}+\phi(\mathbf{u}),\quad T\in[0,20],$ (11)
@@
where $\mathbf{u}=(u(x,y,t),v(x,y,t))^{T}\in\mathbb{R}^{2}$ are two interactive components, $D$ is the diffusion matrix, and $\mathbf{R}(\mathbf{u})$ is source term for the reaction. We fix the parameters to $\alpha=0.01,\beta=1.0,D=\begin{pmatrix}0.05&0\\ 0&0.1\end{pmatrix}$, where the system is known to form Turing patterns [33]. Assuming that we do not know the exact form of the reaction term $\phi(\mathbf{u})$, the resolved part is the diffusion term and the unresolved part is the non-linear reaction term. The ground truth data are calculated using the semi-implicit Crank-Nicolson scheme with full explicit discretization for the non-linear term. As we do not know the reaction term is pointwise in space, we use a convolutional neural network-based model to learn it based on ordinary least squares (OLS). Specifically, we solve the least squares problem $\min_{\theta}\mathbb{E}_{\mathbf{u}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|^{2}$ using Adam [22]. Moreover, we conduct an ablation study where the unresolved model consists of the exact form of the non-linear term as in Eq. 11, plus a random Gaussian noise whose variance is set equal to the optimization error of the OLS estimator, which is $10^{-4}$ in the following experiments. By comparing this and the simulation based on the OLS model, we can conclude whether or not the trajectory error is caused mainly by the estimation error of the unresolved model, or by the shift in distribution of the input data fed to the unresolved model. More details on the data generation and optimization procedure is described in the supplementary materials LABEL:numeric-setup. We simulate this PDE with a test initial condition sampled from the same distribution as training trajectories and compare the results with the ground truth. We present the comparison in Fig. 1.
@@
Figure 1: The first three rows display the snapshots of ground truth, simulated fields, and ablation study simulation at time steps: 0, 100, 200, 300, 500. (d) plots the relative error and distribution shift between two flow configurations at corresponding time. The error is defined as $\frac{\left\|\widehat{\mathbf{u}}_{t}-\mathbf{u}_{t}\right\|_{2}}{\left\|\mathbf{u}_{t}\right\|_{2}}$ at each time step $t$ instead of the whole trajectory as in Eq. 9. The plotted distribution shift roughly measures the average discrepancy between the distributions of $\mathbf{u}_{k}$ and $\widehat{\mathbf{u}}_{k}$ at time step $k$, and is calculated using an autoencoder explained in Section 3.2. Although the OLS estimator and ablation-study estimator share the same error magnitude for unresolved model, their long time trajectory performance form sharp contrast, indicating the optimization error of the unresolved model could not explain the failure in long time trajectory prediction.
@@
The first three rows display the snapshots of ground truth, simulated fields, and ablation study simulation at time steps: 0, 100, 200, 300, 500. (d) plots the relative error and distribution shift between two flow configurations at corresponding time. The error is defined as $\frac{\left\|\widehat{\mathbf{u}}_{t}-\mathbf{u}_{t}\right\|_{2}}{\left\|\mathbf{u}_{t}\right\|_{2}}$ at each time step $t$ instead of the whole trajectory as in Eq. 9. The plotted distribution shift roughly measures the average discrepancy between the distributions of $\mathbf{u}_{k}$ and $\widehat{\mathbf{u}}_{k}$ at time step $k$, and is calculated using an autoencoder explained in Section 3.2. As can be observed, it does not take a very long time for the distribution shift issue to be severe enough using ordinary least squares estimator. Comparing the OLS error and distribution shift with the those of the ablation study, we conclude that the estimation error is not the main cause of the trajectory error, since the ablation-study trajectory has the same magnitude of estimation error, but has much smaller trajectory error and distribution shift. Moreover, in both cases the relative error of the trajectory has the same trend as the distribution shift, indicating correlation.
@@
Indeed, this phenomenon is ubiquitous in modern MLHS [51, 56, 46]. As a second example, we solve the Navier-Stokes equation Eq. 3 using projection method Eq. 4 where pressure solver is replaced by a learned convolutional neural network predictor. The predictor is optimized over data pair of fluid velocity and pressure obtained from high-fidelity simulation to attain an error of $10^{-5}$. Due to the complexity of projection solver and grid issue, we do not conduct ablation study in this case.
@@
Figure 2: Configurations of the Navier-Stokes equation simulation with Reynolds number $200$ at time step: 0, 60, 120, 180, 240: (a) ground truth (b) simulated fluid field using OLS estimator. We only show velocity field over part of the domain. The dynamics blow up in the last configuration, which does not show up in the case of reaction-diffusion equation. This can be explained by the fact that the Navier-Stokes equation is more unstable than the reaction-diffusion equation. In (c), we illustrate the error along the trajectories, the distribution shift.
@@
The distribution shift in Fig. 2 turns out to be more severe when we simulate the incompressible Navier-Stokes equation. After $160$ time steps, the fluid configuration becomes rough and un-physical. Here, we again observe that the DS and error increase together, suggesting correlation. In the next section, we will make this connection precise, and furthermore develop principled methods to control the distribution shift, thereby improving the prediction fidelity.

### v2

@@ 2.3 Learning the unresolved model and distribution shift
We assume that the resolved part $\mathcal{L}$ is known, or we at least have access to a gray box that can perform its evaluation and compute its gradients. While the unresolved part is unknown, we can access data tuples $\left\{(\mathbf{u}_{1},\mathbf{y}_{1},t_{1}),(\mathbf{u}_{2},\mathbf{y}_{2},t_{2}),\cdots,(\mathbf{u}_{N},\mathbf{y}_{N},t_{N})\right\}$ which are obtained either by physical experiments or accurate but expensive numerical simulations. The goal is hence to construct an approximate mapping $(\mathbf{u}_{k},t_{k})\mapsto\mathbf{y}_{k}$ from this dataset. Viewing this as a supervised learning task, there are many existing methods based on empirical risk minimization [18, 19, 27]. One of the simplest methods is to minimize the mean squared error. Given a parameterized model $\phi_{\theta}$, e.g. a Gaussian process or a neural network, we learn from data the optimal value of the parameter $\theta$ that minimizes the error, i.e. $\widehat{\theta}=\arg\min_{\theta}\mathbb{E}_{(\mathbf{u},\mathbf{y})}\left\|\mathbf{y}-\phi_{\theta}(\mathbf{u},t)\right\|_{2}^{2}$. However, due to the dynamical structure, the i.i.d. assumption in statistical learning [19] no longer holds in MLHS. In particular, data belonging to the same dynamical trajectory will have a large correlation. Moreover, an appropriate performance evaluation is not necessarily the precise identification of the system [33] but the accuracy of the prediction of the trajectory over long times. This is also known as a priori and a posteriori performance in the closure modeling literature  [48]. Therefore, it is reasonable to expect that simple empirical risk minimization (which many existing MLHS works employ [54, 45]) may not be the optimal method. We will demonstrate this by using the least squares estimator as a baseline comparison to our approach in Section 4.
@@
After determining the parameters of the unresolved model, one can perform hybrid simulations to obtain new trajectories, i.e.
(7) $\widehat{\mathbf{u}}_{k+1}=\widehat{\mathbf{u}}_{k}+L(\widehat{\mathbf{u}}_{k},\widehat{\mathbf{y}}_{k},t_{k})\Delta t,\quad\widehat{\mathbf{y}}_{k}=\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k},t_{k}).$
@@
Notice we add hat superscript to all the quantities related to this simulated dynamics to distinguish them from the ground truth $(\mathbf{u}_{k},\mathbf{y}_{k})$. We measure the performance of the simulator by the error along the whole trajectory, i.e.
(8) $\min\sum_{k=1}^{n}\left\|\mathbf{u}_{k}-\widehat{\mathbf{u}}_{k}\right\|_{2}^{2},$
@@
where $\mathbf{u}_{k}$s belongs to the true trajectory and $\widehat{\mathbf{u}}_{k}$s the simulated one with the same initial condition.
@@
We begin by illustrating the issue of distribution shift by analyzing the trajectory error. For simplicity of presentation, we assume that the model hypothesis space $\{\phi_{\theta}:\theta\in\Theta\}$ is bias-free, meaning that there exists $\theta^{*}\in\Theta$ such that $\phi=\phi_{\theta^{*}}$. In the general case where the hypothesis space is universal but not closed, this equality would be replaced by an approximate one, but the argument follows analogously. Comparing Eq. 1 and Eq. 7, we have
(9) $\widehat{\mathbf{u}}_{k+1}-\mathbf{u}_{k+1}=$ $\ \widehat{\mathbf{u}}_{k}-\mathbf{u}_{k}+\left(L(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))-L(\mathbf{u}_{k},\phi_{\theta_{*}}(\mathbf{u}_{k}))\right)\Delta t$
$+\left(L(\widehat{\mathbf{u}}_{k},\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k}))-L(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))\right)\Delta t.$
@@
The error introduced by moving one step forward is thus decomposed into two parts: the error associated with the estimator $\widehat{\theta}:L(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))-L(\mathbf{u}_{k},\phi_{\theta_{*}}(\mathbf{u}_{k}))$, and the other error $L(\widehat{\mathbf{u}}_{k},\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k}))-L(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))$. The former resembles the error appearing in classical statistical inference problem under the i.i.d. setting since $\mathbf{u}_{k}$ follows an identical distribution of the training dataset. While the latter is different in the sense that $\widehat{\mathbf{u}}_{k}$ and $\mathbf{u}_{k}$ may come from distributions of their respective driving dynamics, which are different since $\theta_{*}\neq\widehat{\theta}$ due to noisy labels, few data points, or having an under-determined system. The difference between MLHS and statistical learning is also emphasized in [59]. The latter error $L(\widehat{\mathbf{u}}_{k},\phi_{\widehat{\theta}}(\widehat{\mathbf{u}}_{k}))-L(\mathbf{u}_{k},\phi_{\widehat{\theta}}(\mathbf{u}_{k}))$ is rather akin to the issue of stability in numerical partial differential equations: although the error in each step is relatively small, it can accumulate exponentially as time step iterates, driving $\widehat{\mathbf{u}}_{k}$ to completely different regimes where we have not observed in the data $\mathbf{u}_{k},k=1,\cdots,N$. As a result, the data-driven model $\widehat{\phi}$ can no longer be trusted to be accurate, since it is trained on data from a different distribution. This may then lead to a vicious cycle, where further deterioration of trajectory error occurs, leading to an increasing discrepancy between the distributions. Consequently, one can understand the distribution shift as a large discrepancy between the true distribution $\rho$ of $\mathbf{u}_{k}$ and simulated distribution $\widehat{\rho}$ of $\widehat{\mathbf{u}}_{k}$. Concretely, we adopt the following definition for a simulation algorithm Eq. 7 to have distribution shift:
(10) $\limsup_{t\rightarrow\infty}\frac{\log d(\rho_{t},\widehat{\rho}_{t})}{t}\geq c>0,$
where $\rho_{t},\widehat{\rho}_{t}$ denote the distributions of the ground truth $\mathbf{u}(t)$ and the simulation $\widehat{\mathbf{u}}(t)$ starting from the initial condition with same distribution $\rho_{0}$. $d(\cdot,\cdot)$ is an appropriate distance between probability distributions. We emphasize that the definition of the distribution shift relates to the time because of the dynamic nature of the problem, which is different from traditional machine learning settings [25]. The problem of distribution shift occurs commonly in MLHS. For example, [31, 61, 43] all observe unphysical solutions when substituting Reynolds stresses models learned from direct numerical simulation database into RANS. Ref. [61] attributes this to the ill-conditioning of the RANS problems and propose a method to treat Reynolds stresses implicitly. References [50, 49] deal with the subgrid-scale modeling in LES simulations by using neural networks. Improved performance in simulations is achieved by taking into consideration the whole trajectory accuracy in the loss function, which implicitly combats distribution shift. Despite several empirical successes, a general theoretical framework to understand and combat distribution shift in MLHS is lacking. The current paper aims to make progress in this direction.
@@
Let us illustrate the problem of distribution shift using numerical examples. We solve the 2D FitzHugh-Nagumo reaction-diffusion equation over a periodic domain $[0,6.4]\times[0,6.4]$, whose dynamic is given by
(11) $\frac{\partial\mathbf{u}}{\partial t}$ $=D\Delta\mathbf{u}+\phi(\mathbf{u}),\quad T\in[0,20],$
$\phi(\mathbf{u})$ $=\phi(u,v)=\begin{pmatrix}u-u^{3}-v+\alpha\\ \beta(u-v)\end{pmatrix},$
@@
where $\mathbf{u}=(u(x,y,t),v(x,y,t))^{T}\in\mathbb{R}^{2}$ are two interactive components, $D$ is the diffusion matrix, and $\phi(\mathbf{u})$ is source term for the reaction. We fix the parameters to $\alpha=0.01,\beta=1.0,D=\begin{pmatrix}0.05&0\\ 0&0.1\end{pmatrix}$, where the system is known to form Turing patterns [36]. Assuming that we do not know the exact form of the reaction term $\phi(\mathbf{u})$, the resolved part is the diffusion term and the unresolved part is the non-linear reaction term. The ground truth data is calculated using the semi-implicit Crank-Nicolson scheme with fully explicit discretization for the non-linear term. Assuming we do not know that the reaction term is pointwise in space, we use a convolutional neural network-based model to learn it based on ordinary least squares (OLS). Specifically, we solve the least squares problem $\min_{\theta}\mathbb{E}_{\mathbf{u}}\left\|\phi_{\theta}(\mathbf{u})-\phi(\mathbf{u})\right\|_{2}^{2}$ using Adam [24]. Moreover, we conduct an ablation study where the unresolved model consists of the exact form of the non-linear term as in Eq. 11, plus a random Gaussian noise whose variance is set equal to the optimization error of the OLS estimator, which is $10^{-4}$ in the following experiments. By comparing this and the simulation based on the OLS model, we can identify the source of the accumulating trajectory error as either the estimation error of the unresolved model or the shift in the distribution of the input data fed to the unresolved model. More details on the data generation and optimization procedure is described in subsection 2.1 and 2.2 of supplementary materials . We simulate this PDE with a test initial condition sampled from the same distribution as training trajectories and compare the results with the ground truth. We present the comparison in Fig. 1.
@@
Figure 1: The first three rows display the snapshots of ground truth, simulated fields, and ablation study simulation at time steps: 0, 100, 200, 300, and 500. (d) plots the relative error (solid line) and distribution shift (dash line) between two flow configurations at the corresponding time. The error is defined as $\frac{\left\|\widehat{\mathbf{u}}_{t}-\mathbf{u}_{t}\right\|_{2}}{\left\|\mathbf{u}_{t}\right\|_{2}}$ at each time step $t$ instead of the whole trajectory as in Eq. 8. The plotted distribution shift roughly measures the average discrepancy between the distributions of $\mathbf{u}_{k}$ and $\widehat{\mathbf{u}}_{k}$ at time step $k$, and is calculated using an autoencoder explained in Section 3.2. Although the OLS estimator and ablation-study estimator share the same error magnitude for unresolved models, their long-time trajectory performance forms sharp contrasts, indicating the optimization error of the unresolved model could not explain the failure in long-time trajectory prediction.
@@
The first three rows display the snapshots of ground truth, simulated fields, and ablation study simulation at time steps: 0, 100, 200, 300, and 500. (d) plots the relative error (solid curve) and distribution shift (dotted curve) between two flow configurations at the corresponding time. The error is defined as $\frac{\left\|\widehat{\mathbf{u}}_{t}-\mathbf{u}_{t}\right\|_{2}}{\left\|\mathbf{u}_{t}\right\|_{2}}$ at each time step $t$ instead of the whole trajectory as in Eq. 8. The plotted distribution shift roughly measures the average discrepancy between the distributions of $\mathbf{u}_{k}$ and $\widehat{\mathbf{u}}_{k}$ at time step $k$, and is calculated using an autoencoder explained in Section 3.2. As can be observed, it does not take a very long time for the distribution shift issue to be severe enough using the OLS estimator. Comparing the OLS error and distribution shift with those of the ablation study, we conclude that the estimation error is not the main cause of the trajectory error, since the ablation-study trajectory has the same magnitude of estimation error, but has a much smaller trajectory error and distribution shift. Moreover, in both cases the relative error of the trajectory has the same trend as the distribution shift, indicating correlations.
@@
Indeed, this phenomenon is ubiquitous in modern MLHS [59, 64, 54]. As a second example, we solve the Navier-Stokes equation Eq. 3 using projection method Eq. 4 where the pressure solver is replaced by a learned convolutional neural network predictor. The predictor is optimized over data pairs of fluid velocity and pressure obtained from high-fidelity simulation to attain an error of $10^{-5}$. Due to the complexity of the projection solver and grid issue, we do not conduct an ablation study in this case.
@@
Figure 2: Configurations of the Navier-Stokes equation simulation with Reynolds number $200$ at time step: 0, 60, 120, 180, 240: (a) ground truth (b) simulated fluid field using OLS estimator. We only show the velocity field over part of the domain. The dynamic blows up in the last configuration, which does not show up in the case of reaction-diffusion equations. This can be explained by the fact that the Navier-Stokes equation is more unstable than reaction-diffusion equations. In (c), we illustrate the error along the trajectories, the distribution shift.
@@
The distribution shift in Fig. 2 turns out to be more severe when we simulate the incompressible Navier-Stokes equation. After $160$ time steps, the fluid configuration becomes rough and un-physical. Here, we again observe that the DS and error increase together, suggesting correlations. In the next section, we will make this connection precise, and develop principled methods to control the distribution shift, thereby improving the prediction fidelity.